OpenStudy (anonymous):

Two vector A=(3, 4, -5) and B=(-1, 1, 2). Determine the magnitude of A+B and the direction (angle) of A+B ?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

did you find \(A+B\)?

OpenStudy (anonymous):

Yes, it is \[\sqrt{38}\] and the angle I found is 121 but the answer it the book says it is 71 =(

OpenStudy (misty1212):

oh how did you find the angle (i actually am not sure what angle means in three dimensions)

OpenStudy (anonymous):

A.B( dot product) = \[\left| A \right|.\left| B \right|\] .cos

OpenStudy (misty1212):

doesn't that find the angle between the two vectors?

OpenStudy (anonymous):

Yes, think the book has the wrong answer! thank you anyway ^^